# Probabilistically-sound beam search with masked language models

### v1

@@ 6.4 Ablations
Ablation 1 dramatically worsens the performance of HCB beam search, as we show in Table 7. Ablation 2 also deteriorates the performance of HCB beam search, although by less than that of ablation 1.
We conclude that $p([\operatorname{M}]|\mathbf{x})$ is sensitive to both the context $\mathbf{x}$ and input token $[\operatorname{M}]$, and we hypothesize that this contributes to the success of HCB beam search. Since results are far worse when we randomly replace $\mathbf{x}$ than when we randomly replace $[\operatorname{M}]$, we hypothesize that the context $\mathbf{x}$ is a more important factor to the success of HCB beam search.
@@ 7 Discussion
In Section 6, we see that HCB beam search achieves comparable performance to standard beam search across languages and that superiority of either method is usually model-dependent, rather than data-dependent: for instance, HCB beam search consistently outperforms standard beam search with BERT-base, but consistently underperforms with RoBERTa, across datasets, as shown by Table 1.
@@
For this reason, we might expect standard beam search to function well exactly when this conditional independence holds. BERT’s training procedure is known to encourage this conditional independence Devlin et al. (2019), and so we might expect independence to be a more reasonable assumption for more finely optimized BERT models, such as RoBERTa. Indeed, this hypothesis is consistent with Table 1, which finds that standard beam search outperforms HCB beam search over all datasets considered when RoBERTa is used, but not when BERT-base is used. We hypothesize this may also explain why standard beam search outperforms HCB beam search for ESM2.
@@
$p(x_{j}|\mathbf{x}_{<j},\mathbf{[\operatorname{M}]}_{>j})=p(x_{j}|\mathbf{x}_{<j});$
@@ 8 Conclusion
In this work, we develop and apply theoretically-sound methods to use pretrained MLMs for text infilling, a task with important applications spanning a wide variety of domains. We clarify the conditions under which it is theoretically sound to perform text infilling with MLMs using standard beam search. For instances when these conditions do not hold, we introduce HCB beam search as a probabilistically-justified modification with no additional computational complexity and demonstrate its superiority to standard beam search in the expected conditions. Future work exploring HCB beam search in the context of other MLMs and in other domains can help further elucidate the contexts in which HCB beam search is beneficial, ultimately facilitating the adaptation of MLMs for text infilling and related tasks in low-data or compute settings.
@@ 9 Limitations
One limitation is that we do not explicitly know whether conditional independence assumptions hold or fail for a given model, despite hypothesized heuristics. Another limitation is the lack of intuition regarding the use of the mask token as a pivot – future work should more rigorously investigate optimal pivot choices and the interpretability of these entities. Additionally, a further understanding of the objective and training conditions under which an MLM satisfies conditional independence assumptions is also needed. Finally, the proposed HCB beam search relies on the presence of some token in a given model’s output distribution which flexibly behaves as in-distribution; while some MLMs maybe include special characters in their output layers, others may not. A potential risk of our method is its ability to use MLMs outside of their original contexts, which could have unforeseen consequences and lead to unpredictable behavior.
@@ Appendix A Proofs
$P(x^{i}_{j}=x|\mathbf{y}^{i}_{<j}=\mathbf{z}_{<j},\mathbf{y}^{i}_{>j}=\mathbf{[\operatorname{M}]}_{>j})$
for arbitrary $x$ and $\mathbf{z}$. As the event $\{\mathbf{y}^{i}_{k}=[\operatorname{M}]\}$ occurs if and only if $Z_{ik}=1$, and thus with probability $p$ independently of $\mathbf{x}$ for all $i\in[m],j\in[n]$, the conditional independence claim follows. ∎
@@
for all $n$, then we recover the conditional independence assumption by induction on $n$. ∎
Appendix B Additional figures
@@
Model BERT-base BERT-large DistilBERT RoBERTa
HCB Left-to-Right 10.86 11.92 6.12 24.74
HCB Best-to-Worst 11.43 12.58 6.44 25.29
Ablation 1 Left-to-Right 4.10 3.83 1.86 9.42
Ablation 1 Best-to-Worst 4.63 4.08 2.06 9.86
Ablation 2 Left-to-Right 8.94 10.65 4.82 16.34
Ablation 2 Best-to-Worst 9.71 11.16 5.23 17.61
Table 7: Top-1 accuracy comparison of HCB beam search to ablations 1 and 2, across models, when infilling gaps consisting of a (uniformly random) number of tokens between 2 and 5.
Model BERT-base BERT-large DistilBERT RoBERTa
HCB Left-to-Right 10.86 11.92 6.12 24.74
HCB Best-to-Worst 11.43 12.58 6.44 25.29
Standard Left-to-Right 9.70 11.31 5.87 24.98
Standard Best-to-Worst 10.63 12.32 6.39 25.81
Pure sampling 8.09 9.57 4.00 20.72
Sampling w/ $T=0.25$ 9.41 10.97 5.66 24.59
Sampling w/ $T=0.5$ 9.25 10.74 5.40 24.03
Sampling w/ $T=0.75$ 8.79 10.31 4.85 22.89
Nucleus sampling, $p=0.9$ 8.45 9.99 4.30 22.57
Ablation 1 (Context Scramble) 4.63 4.08 2.06 9.86
Ablation 2 (Random Token Swap) 9.71 11.16 5.23 17.61
Table 8: Top-1 accuracy at infilling a random number (uniform between 2 and 5) of missing tokens on 100K examples from Brown corpus, across models, with beam size 5. For ablations, we show only best-to-worst results since they strictly outperformed left-to-right results.
Figure 5: Comparison of HCB beam search with standard beam search. Evaluated on 10,000 examples each from the SNLI and XSUM datasets.
@@
Figure 7: Average $p([\operatorname{M}]|\cdot)$ for fixed random masks across 50 test examples over the course of BERT training.

### v3

@@ 6.4 Ablations
Ablation 1 dramatically worsens the performance of HCB beam search, as we show in Table 3. Ablation 2 also deteriorates the performance of HCB beam search, although less so than Ablation 1. We conclude that $p([\operatorname{M}]|\mathbf{x})$ is sensitive to both the context $\mathbf{x}$ and input token $[\operatorname{M}]$, and we hypothesize that this contributes to the success of HCB beam search. Since results are far worse when we randomly replace $\mathbf{x}$ than when we randomly replace $[\operatorname{M}]$, we hypothesize that the context $\mathbf{x}$ is a more important factor to the success of HCB beam search.
@@ 7 Discussion
In Section 6, we see that HCB beam search achieves comparable performance to standard beam search across languages and that superiority of either method is usually model-dependent, rather than data-dependent: for instance, HCB beam search consistently outperforms standard beam search with BERT-base, but consistently underperforms with RoBERTa across datasets as shown by Table 2.
@@
$p(x_{j}|\mathbf{x}_{:j},\mathbf{[\operatorname{M}]}_{j+1:})=p(x_{j}|\mathbf{x}_{:j});$
@@
To explore this hypothesis, we trained two BERT models from scratch under different training conditions which might affect the validity of the conditional independence assumption. Specifically, one model was trained with dynamic masking, where the data is re-masked every epoch, and the other with static masking, where the data is masked only once at the start of training. After assessing a large number of checkpoints across both models, we do not observe HCB beam search becoming significantly less effective compared to standard beam search as training progressed. As performance appears to be model-specific, we recommend that researchers using beam search evaluate HCB beam search in their setting, as it requires minimal modification from standard beam search.
@@ 8 Conclusion
In this work, we develop and apply theoretically-sound methods to use pretrained MLMs for text infilling, a task with important applications spanning a wide variety of domains. We clarify the conditions under which it is theoretically sound to perform text infilling with MLMs using standard beam search. We introduce HCB beam search as a probabilistically-justified modification with no additional computational complexity and demonstrate its superiority to standard beam search with a suite of models. Future work exploring HCB beam search can help further elucidate the contexts in which HCB beam search is beneficial, especially in low-data or compute settings.
@@ 9 Limitations
One limitation is that we do not explicitly know whether conditional independence assumptions hold for a given model, despite hypothesized heuristics. Another is the lack of intuition regarding the use of [mask] as a pivot – future work should more rigorously investigate optimal pivot choices and the interpretability of these entities.
Moreover, the proposed HCB beam search relies on the presence of some token in a given model’s output distribution which flexibly behaves as in-distribution; while some MLMs maybe include special characters in their output layers, others may not. A potential risk of our method is its ability to use MLMs outside of their original contexts, which could have unforeseen consequences and unpredictable behavior.
Finally, our proposed method only applies to settings with a fixed number of missing tokens as opposed to the more general setting addressed by others (Shen et al., 2020). Nonetheless, we find this is a realistic assumption in some important applications, such as ancient text restoration and protein engineering. In the case of damaged inscriptions, domain experts posit an estimated number of missing characters based on physical distance (Bruun and Edmondson, 2014). In protein engineering, a reasonable constraint is maintaining a fixed protein sequence length while only modifying amino acids in functional regions. In both areas, the fixed length assumption has been used in cutting-edge models (Assael et al., 2022; Wang et al., 2022).
@@ Appendix A Proofs
$p(x^{i}_{j}=x|\mathbf{y}^{i}_{:j}=\mathbf{z}_{:j},\mathbf{y}^{i}_{j+1:}=\mathbf{[\operatorname{M}]}_{j+1:})$
for arbitrary $x$ and $\mathbf{z}$. As the event $\{\mathbf{y}^{i}_{k}=[\operatorname{M}]\}$ occurs if and only if $Z_{ik}=1$, and thus with probability $p$ independently of $\mathbf{x}$ for all $i\in[m],j\in[n]$, the conditional independence claim
$p(x^{i}_{j}|\mathbf{x}^{i}_{:j},\mathbf{[M]}_{j+1:})=p(x^{i}_{j}|\mathbf{x}^{i}_{:j})$
follows. ∎
@@
for all $n$, then we recover the conditional independence law
$p(x_{j}|\mathbf{x}_{:j},[\operatorname{M}]_{j+1:})=p(x_{j}|\mathbf{x}_{:j})$
by induction on $n$. ∎
Appendix B Algorithms
Initialize $S=\{(0,\emptyset)\}$
for $i\in\{1,\ldots,n\}$ do
     for $(\ell,(x_{1},\ldots,x_{i-1}))\in S$ do
         $f(\cdot)\leftarrow\log p(\cdot|{x_{1},\ldots,x_{i-1}})$ (1 forward pass.)
         Append to $S$: $(\ell+f(x),(x_{1},\ldots,x))$ for every $x$.
     end for
[…]
end for
return S
Appendix C Additional figures
@@
Figure 4: Performance of HCB beam search versus standard beam search methods across three English MLMs and DistilBERT, with beam size 20.
@@
[…]